OpenStudy (anonymous):

Ginny's class ordered 4 pizzas. They ate seven-twelfths of each pizza. How many pizzas did they eat? 2 and one-third 3 and one-fourth 2 and one-fourth 1 and three-fourths

OpenStudy (anonymous):

help

OpenStudy (tkhunny):

\(\dfrac{7}{12}\cdot 4 = 7\cdot\dfrac{4}{12}\) Are we getting anywhere?

OpenStudy (anonymous):

"Of" means multipy. So 7/12 of 4 translates to\[\frac{7}{12}(4)=\frac{7}{3}=2\frac{1}{3}\]